# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v6

@@ Abstract
Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence. Existing evaluation frameworks, which are centered on language or perception tasks, fail to capture generality at its core and offer no guidance. The Artificial General Intelligence Testbed (AGITB) is a novel, freely available benchmarking suite consisting of thirteen core requirements, twelve of which are implemented as fully automatable tests designed to assess low-level cognitive precursors through binary signal prediction. AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing. Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles. While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence. A reference implementation of AGITB is available on GitHub.
Keywords: artificial general intelligence, benchmarking, generalization, symbol grounding problem, temporal sequence prediction
@@ 1 Introduction
Despite rapid advancements in machine learning and neural network architectures, artificial intelligence (AI) systems still fall short of flexible, robust human intelligence. Marcus and Davis (2020) observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. Mitchell (2025) further summarized why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted. The surface-level competence obscures a deeper lack of understanding—an essential prerequisite for artificial general intelligence (AGI).
The progress in AI has led to growing speculation that AGI is near. However, such claims remain difficult to substantiate without a rigorous and informative metric. Assessing progress toward AGI requires more than specialized benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing (1950). However, no existing metric fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability. They typically assess surface-level competence rather than underlying mechanisms of generalization. In response to these limitations, this paper introduces the artificial general intelligence testbed (AGITB), a novel framework for evaluating foundational cognitive abilities in artificial intelligence systems.
AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising thirteen core requirements—twelve of which are implemented as fully automatable, domain-agnostic tests. While AGITB does not aim to assess consciousness or semantic comprehension, it provides a principled mechanism for distinguishing narrow AI systems from those exhibiting generalizable, adaptive behavior. A comparative analysis with the Abstraction and reasoning corpus (ARC) introduced by Chollet (2019) is included to contextualize AGITB within the broader landscape of AGI evaluation.
@@ 2 Background
The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities. However, despite these advances, a persistent reluctance remains to attribute “intelligence” to machines. This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits—such as consciousness, self-awareness, and subjective experience—which remain elusive in artificial systems.
Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have often shifted. For instance, the success of Deep Blue against Garry Kasparov in chess (a task formerly seen as a benchmark for AGI) was quickly reframed as a triumph of brute-force computation rather than genuine intelligence. Such redefinitions risk obscuring real milestones. As AI approaches human-level capabilities in certain areas, we may inadvertently set a perpetually receding goalpost for AGI, failing to recognize it even when achieved.
Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics. This is partly due to the absence of a universally accepted AGI benchmark. Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified. However, such specialized benchmarks favor narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence. Ironically, some of these benchmarks are now so specialized that humans have difficulty with them.
@@ 2.1 A Benchmark That Only Humans and AGI Can Meet
An effective AGI test should be intuitively solvable by humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching. Such a test must demand capabilities that transcend memorization or domain-specific heuristics, requiring generalization, abstraction, and adaptive reasoning.
A valid AGI evaluation must either (1) reveal a fundamental cognitive gap between humans and machines or (2) define a behavioral capacity that current artificial systems cannot replicate. The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets. The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass (1997).
In alignment with this biologically grounded perspective, the artificial general intelligence testbed departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level. While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad (1990).
AGITB thus adopts a more elemental approach. Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalize patterns in raw binary signals independently of semantics or prior training. Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting. This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory data, not disembodied symbols.
@@ 3 Artificial General Intelligence Testbed
The primary objective of the artificial general intelligence testbed is to support the development and evaluation of AGI by defining a clear set of requirements that a model must meet to qualify as generally intelligent. A model must satisfy all requirements in order to validly claim success on the benchmark.
@@ 3.1 Components
AGITB requires the user to define two components: the cortex, representing the AGI system under evaluation, and the input, which encodes the data samples delivered to the cortex over time. The cortex is treated as a black box that predicts future inputs based on the observed history of prior signals.
Each input consists of binary-encoded data from (virtual) sensors and actuators. It is composed of a fixed number of bits, with each bit corresponding to a distinct input channel, such as an individual pixel, audio band, or actuator feedback signal. An input represents a snapshot of multiple parallel signals at a single point in time, as illustrated in Figure 1a. While spatial information is encoded within the structure of each input, semantic richness emerges through the temporal unfolding of input sequences, as shown in Figure 1b. Although the spatial and temporal dimensions are orthogonal, their interaction produces structured patterns to which the cortex must adapt.
Figure 1: (a) Example of a $2\times 3$-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
(b) Example of an input sequence illustrating a temporal pattern with a period of 4.
@@ 3.2 Operation
The testbed presents the cortex with a stream of inputs over time. At each time step $t$, the cortex receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions. In cases where predictions are not exact, proximity to the correct value is considered, which indicates the model’s capacity for generalization and adaptation.
Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps. At the previous step (t1t-1), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data. Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
AGITB asserts expectations about the resulting state and behavioral dynamics of cortex models in specific scenarios using randomly generated test inputs. Rather than relying on arbitrary score thresholds, AGITB evaluates each model by comparing it against itself under controlled conditions. Since the internal state of the cortex is not accessible to the testbed for direct comparison, the user must define a custom criterion for determining model equality.
To execute AGITB, users must specify one parameter: the temporal pattern length. This value determines the period of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to. Longer periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities. However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity. As such, the temporal pattern length should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
3.3 The 13 Essential Requirements
A reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
Requirement 1 – Genesis
@@ 3.3 The 13 Essential Requirements
AGITB rests on the foundational assumption that general-purpose learning systems—such as the brain—do not possess an innate understanding of external inputs. Instead, they acquire understanding through experiential interaction with their environment. Each system must independently construct meaning from raw sensory data without relying on pre-encoded semantics. While some reflexes may be genetically predetermined, these do not constitute genuine understanding. In this respect, all cortical systems begin from an unbiased initial state, shaped entirely by the input received over time.
A model that predicts anything other than an empty (spike-free) input before receiving any data is considered biased, as it encodes implicit assumptions about an unwarranted future. To maintain neutrality, such models must initialize to predict only empty patterns and remain unbiased until influenced by actual input.
Two cortex instances that have received no input should compare as equal, as neither has been influenced by prior experience. Here, “empty” refers to a lack of informational content, not structural capability. The cortex must possess an intrinsic organizational architecture capable of learning, even in the absence of prior data.
@@
Every input modifies the cortex’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
@@
Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function. Although small stochastic effects may occur, they do not undermine the structured, rule-based nature of neural processing. Similarly, in AGITB, two functionally identical cortex models must have experienced identical input histories. Any difference in input must lead to a divergence in state, reinforcing the principle that behavior and internal state are fully determined by input alone. The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave (2016) in his discussion of free will.
@@
Assertion: Changing the input order results in a different cortex state.
Because the cortex updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation. The ability to recognize and exploit temporal structure is therefore a defining property of intelligent systems.
@@
Assertion: The cortex must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
Assertion: The cortex cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
@@
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner and Kistler (2002) showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning. Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
Requirement 7 – Adaptability
Assertion: The cortex must be able to adapt to a temporal pattern with the user-specified period.
Assertion: The cortex must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
This test evaluates the model’s capacity to learn across temporal scales. Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity. The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
@@
Assertion: There exists a limit beyond which the cortex can no longer adapt, even to patterns that would otherwise be learnable.
Cognitive systems inevitably reach adaptation limits as their finite resources become saturated. This test assesses whether such limits emerge over time when the cortex is exposed to input sequences that would otherwise be learnable.
@@
The structure and regularity of an input sequence affect the rate at which a model adapts, where adaptation time is defined as the number of iterations required to correctly predict the entire temporal pattern. Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
@@
Assertion: Adaptation time depends on the state of the cortex.
Past learning influences how new information is integrated. A cortex with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganization.
@@
Assertion: Distinct cortices may exhibit the same observable behavior.
@@
Identical external outputs can arise from different internal states. This many-to-one mapping underscores that observable behavior alone cannot reveal the underlying structure or history of a model’s internal dynamics.
Requirement 12 – Generalization
Assertion: On average, adapted models achieve higher predictive accuracy than unadapted models after input disruption.
Assertion: On average, adapted models achieve higher predictive accuracy than random guessing after input disruption.
@@
Only models capable of generalization will derive lasting benefit from prior learning when inputs are disrupted. Upon re-exposure to familiar stimuli, such models should, on average, outperform both unadapted models and random baselines in predictive tasks. Average performance is the evaluation criterion because even unadapted or random models may occasionally produce correct predictions by chance, without reflecting genuine understanding or learned structure.
Requirement 13 – Bounded Prediction Latency
@@
A cortex model must provide low-latency signal-level prediction, which is essential for functional viability. Biological brains achieve approximately constant reaction times through massive parallelism, such that per-input latency remains bounded and stable under expected workloads.
@@
Unfortunately, no universal automatic test on serial hardware can verify this property across all possible cortex implementations within reasonable time. Instead, model authors must ensure that their architecture admits efficient parallel execution with bounded latency. In this respect, artificial neural networks—with their naturally parallelizable structure—offer a practical advantage.
3.4 Memorization

### v7

@@ Abstract
Current artificial intelligence systems continue to fall short of human-like general intelligence. Existing evaluation frameworks, which focus on language or perception tasks, fail to capture the essence of generality or provide actionable guidance. The Artificial General Intelligence Testbed (AGITB) introduces a novel benchmarking suite comprising fourteen elementary tests, thirteen of which are fully automated.
AGITB evaluates models on their ability to forecast the next input in a temporal sequence, step by step, without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalisation—that parallel principles of biological information processing. Designed to resist brute-force or memorisation-based strategies, AGITB enforces unbiased and autonomous learning. The human cortex satisfies all tests, whereas no current AI system meets the full AGITB criteria, demonstrating its value as a rigorous, interpretable, and actionable benchmark for evaluating progress toward artificial general intelligence. A reference implementation of AGITB is freely available on GitHub.
Keywords: artificial general intelligence, benchmarking, generalisation, symbol grounding problem, temporal sequence prediction
@@ 1 Introduction
Despite rapid advances in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence. Marcus2020 observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. Mitchell2025 further summarised why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted. Surface-level competence obscures a deeper lack of understanding, which is an essential prerequisite for artificial general intelligence (AGI).
The progress in AI has led to a growing speculation that AGI is near. However, such claims remain difficult to substantiate without a rigorous and informative metric. Assessing progress toward AGI requires more than specialised metrics or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing1950. However, no existing test fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability. The metrics typically assess surface-level competence rather than underlying mechanisms of generalisation. In response to these limitations, this paper introduces the Artificial General Intelligence Testbed (AGITB), a novel benchmark for validating foundational cognitive abilities in artificial intelligence systems.
AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, thirteen of which are implemented as fully automatable and domain-agnostic tests. Although AGITB does not aim to evaluate consciousness or semantic comprehension, it provides a principled framework for distinguishing narrow AI systems from those exhibiting generalisable, adaptive behaviour. To position AGITB within the broader landscape of AGI evaluation, we include a comparative analysis with the Abstraction and Reasoning Corpus (ARC; Chollet2019) and the NeuroBench framework (Yik2025).
@@ 2 Background
The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities. However, despite these advances, a persistent reluctance remains to attribute "intelligence" to machines. This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits—such as consciousness, self-awareness, and subjective experience—which remain elusive in artificial systems.
Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have undergone significant shifts. For instance, the success of Deep Blue against Garry Kasparov in chess (a task formerly seen as a benchmark for AGI) was quickly reframed as a triumph of brute-force computation rather than genuine intelligence. Such redefinitions risk obscuring real milestones. As AI approaches human-level capabilities, we may inadvertently set a perpetually receding goalpost for AGI, failing to recognise it even when achieved.
Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics. This is partly due to the absence of a universally accepted AGI benchmark. Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified. However, such specialised benchmarks favour narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence. Ironically, some of these benchmarks are now so specialised that humans have difficulty with them.
@@ 2.1 A Benchmark That Only Humans and AGI Can Meet
An effective AGI test must be trivial to solve for humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching. Such a test must demand capabilities that transcend memorisation or domain-specific heuristics, requiring generalisation, abstraction, and adaptive reasoning.
A valid AGI evaluation must either (1) reveal and exploit a fundamental cognitive gap between humans and machines or (2) define a behavioural capacity that current artificial systems cannot replicate. The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets. The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass1997.
In alignment with this biologically grounded perspective, the AGITB departs from symbolic, language-based evaluations and instead assesses intelligence at the lowest, signal-processing level. While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad1990.
AGITB thus adopts a more elemental approach. Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalise patterns in raw binary signals. A neural spike by itself contains the smallest amount of information possible and is, as such, grounded but free of other semantics. A binary signal accurately represents the neural spike.
Rooted in the view—advanced by Hawkins2004—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting. This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory spike trains, rather than disembodied symbols.
@@ 3 Artificial General Intelligence Testbed
The testbed supports the development and evaluation of AGI by defining a clear set of requirements that a model under evaluation must meet to qualify as generally intelligent. A model must satisfy all requirements in order to claim success on the benchmark.
The central premise is that the capacity to pass the AGITB constitutes a necessary condition for the existence of AGI. Non-AGI systems cannot meet AGITB’s requirements. Such a premise cannot be proven; it can only be rejected by any non-AGI system that meets the AGITB criteria. The fact that humans can pass the AGITB suggests that AGITB requires a solution beyond the current state-of-the-art AI.
Unfortunately, without a clear definition of intelligence, we can only speculate that AGITB defines the necessary (but not sufficient) building blocks for the emergence of genuine, domain-independent intelligence. This assumption of necessity stems from the fact that, except for human brains, no existing artificial system has met its requirements.
However, an AGITB-passing system is not automatically also a fully fledged AGI; after all, it is not required to express higher-level skills, like reasoning or natural language fluency. However, a failure to meet the AGITB’s requirements likely indicates a deficiency in foundational general cognitive capabilities.
@@ 3.1 Components
AGITB evaluates the user-supplied AGI model as a black box that predicts future inputs from the historical sequence of observed signals. Each input comprises ten bits, the specific semantics of which are immaterial; each bit may correspond to an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
Each input represents a snapshot of multiple parallel signals at a single time step (Figure 1a). Spatial organisation within each input encodes local structure, whereas semantic richness arises from the temporal evolution of the input sequence (Figure 1b). The interaction between spatial and temporal dimensions gives rise to structured patterns that are challenging for the model to adapt to.
Figure 1: (a) Example of a 10-bit input $x$ with 4 bits set.
(b) Example of an input sequence showing a temporal pattern of period four.
@@ 3.2 Operation
The testbed presents the model with a stream of inputs over time. At each time step $t$, the model receives an input $p_{t}$ and is required to predict the subsequent input $p_{t+1}$, as shown in Figure 2. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
Figure 2: Iterative adaptation in discrete time. At the previous time step (t1t-1), the model issued a prediction $x_{t}^{*}$, with the second bit incorrect in this instance. After observing the realized input $x_{t}$, the model updates and subsequently produces a correct one-step-ahead prediction $x_{t+1}^{*}$ for the sequence shown in Figure 1b.
3.2.1 Design
A distinguishing architectural feature of AGITB is its deliberate avoidance of conventional correctness metrics such as accuracy or mean-squared error. The limitations of these metrics are twofold. First, they cannot reliably distinguish between AGI and human performance, or that of non-AGI systems, as contemporary AI models can already surpass humans on standard benchmarks. Second, in an era of elevated expectations driven by specialised generative AI systems, the anticipated AGI performance on such metrics is often set unrealistically high. Even well-educated humans possessing fully developed and highly parallelised brains may underperform relative to current AI systems. It is therefore unreasonable to expect a first-generation AGI, potentially operating on a simplified and computationally constrained simulation of the brain, to match or exceed human-level results. Currently, our understanding of neural mechanisms is insufficient to justify such expectations.
Without conventional metrics or predefined performance thresholds, it can be difficult to determine whether a model has achieved a meaningful level of competence. AGITB addresses this challenge by employing a self-referential evaluation approach, in which the model under test is compared against its own independent instances. Each test constructs a controlled scenario involving one or more such models, whose internal states and behaviours are analysed comparatively. Requiring each test to be passed 100 times renders AGITB an extreme form of stress testing, ensuring that successful performance reflects genuine robustness rather than chance. Success is thus defined in terms of the relative consistency or superiority of model responses, rather than by any external quantitative metric.
For these reasons, and to minimise type I errors, AGITB employs an all-or-nothing criterion: the system under evaluation must successfully pass all fourteen tests. This design choice is justified, as individual tests are solvable by non-AGI systems, whereas the simultaneous satisfaction of all fourteen requirements is likely to demand capabilities beyond narrow intelligence. AGITB therefore posits that what non-AGI systems lack is the property referred to as "intelligence," understood as the synergistic integration of all fourteen tests.
3.3 The 14 AGI Requirements
The reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
Requirement 1 – Unbiased start
@@ 3.3 The 14 AGI Requirements
AGITB rests on the foundational assumption that general-purpose learning systems—such as the brain—do not possess an innate understanding of external inputs. Instead, they learn through interaction with their environment. Each system must independently construct meaning from raw sensory data without relying on pre-encoded semantics. While some reflexes may be genetically predetermined, these do not constitute genuine understanding. In this respect, all cortical systems begin from an unbiased initial state, shaped entirely by the input received over time.
A model that predicts non-empty (non-zero) inputs prior to receiving data is considered biased, as it encodes unwarranted assumptions about the future. To ensure neutrality, models must initialise with empty predictions and remain unbiased until conditioned on actual input.
Two model instances that have received no input should compare as equal, as neither has been influenced by prior experience. Here, "empty" refers to a lack of informational content, not structural capability. The model must possess an intrinsic organisational architecture capable of learning, even in the absence of prior data.
The test operationalises, in empirical form, the philosophical challenge posed by classical symbolic systems: the emergence of meaning without prior grounding.
@@
Every input modifies the model’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
@@
Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function. Although small stochastic effects may occur, they do not undermine the structured, rule-based nature of neural processing. Similarly, in AGITB, two functionally identical models must have experienced identical input histories. Any difference in input must lead to a divergence in state, reinforcing the principle that behaviour and internal state are fully determined by input alone. The brain’s actions are effectively deterministic, though often perceived otherwise due to their immense complexity, as noted by Cave2016 in his discussion of free will.
@@
Assertion: Changing the input order results in a different model state.
Because the model updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation. The ability to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
@@
Assertion: The model must be able to adapt to any minimal-period input sequence that respects proper refractory periods.
Assertion: The model cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
@@
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner2002 showed that input sequences that lack sufficient variability—such as constant or overly repetitive spiking—fail to support effective learning. Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
The refractory period requirement does not preclude the possibility of more general solutions. A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot simulate the absence of such dynamics.
Requirement 7 – Temporal adaptability
Assertion: The model must be able to adapt to a temporal pattern with the user-specified period.
Assertion: The model must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
This test assesses the model’s ability to learn across different temporal scales. Unlike rigid temporal pattern-matching systems, intelligent models should be able to detect and respond to recurring structures, regardless of their exact periodicity. The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
@@
Assertion: There exists a limit beyond which the model can no longer adapt, even to patterns that would otherwise be learnable.
Cognitive systems inevitably reach adaptation limits as their finite resources become saturated. This test assesses whether such limits emerge over time when the model is exposed to input sequences that would otherwise be learnable.
@@
The structure and regularity of an input sequence influence the rate at which a model adapts, where adaptation time is defined as the number of iterations required to predict the entire temporal pattern accurately. Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
@@
Assertion: Adaptation time depends on the state of the model.
Past learning influences how new information is integrated. A model with a well-structured internal state may adapt quickly to familiar or related patterns, whereas unstructured or conflicting states may require more extensive reorganisation.
@@
Assertion: Distinct cortices may exhibit the same observable behaviour.
@@
Identical external outputs can arise from different internal states. This many-to-one mapping underscores that observable behaviour alone cannot reveal the underlying structure or history of a model’s internal dynamics.
Requirement 12 – Denoising
Assertion: A model can recall a sequence despite perturbations.
An intelligent model should be capable of efficiently memorising input patterns in a manner that enables the recall of previously observed sequences even when these are perturbed by noise. When re-exposed to familiar stimuli, such a model is expected, on average, to outperform random guessing in predictive tasks. Average performance is used as the evaluation criterion, since random models may occasionally produce correct predictions by chance without reflecting genuine learning or structural understanding.
Test 12 is designed as a deliberately conservative benchmark. The model is required to outperform random guessing in all 100 independent trials. This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it serves as a stringent robustness check. The design ensures that only large and systematic performance gains result in a pass. Modest or marginal improvements, while potentially real, are intentionally treated as failures, as the objective is to identify only clear and substantial advances in model capability.
Requirement 13 — Generalization
Assertion: The model performs above chance on previously unseen inputs.
@@
Only models capable of generalisation can derive lasting benefit from prior learning. After exposure to a given set of stimuli, such models are expected, on average, to outperform random baselines when predicting future inputs. Average performance is used as the evaluation criterion, since even a random model may occasionally produce correct predictions by chance without demonstrating genuine understanding or the acquisition of structured knowledge.
The generalisation assessment follows the same conservative design as Test 12. By requiring success across all repetitions, the procedure establishes a stringent threshold whereby a single underperforming trial constitutes failure. Consequently, the test exhibits minimal tolerance for noise or marginal effects. This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models demonstrating a clear and consistent advantage achieve a passing result.
Requirement 14 – Bounded Prediction Latency
@@
A model must provide low-latency, signal-level predictions to ensure functional viability. Biological brains achieve approximately constant reaction times through extensive parallel processing, such that per-input latency remains bounded and stable under expected workloads.
@@
No universal automated test executed on serial hardware can feasibly verify this property across all possible model implementations. Consequently, model developers are responsible for ensuring that their architectures permit efficient parallel execution with bounded latency. In this regard, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
3.4 Search space